MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2008/04/12
OBSERVER: Fukui Paul
CONDITION: clear
COMMENT: D3566(300sec), D3569(30sec), D3570(60sec) are OK. We observed gb2 as the same frequency as gb5 and 9 for MOA-75.
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 19:15 END 06:23
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 12477-12515 | 39 | 246 | 14760 |
LMC | 2705-2710 | 6 | 34 | 10200 |
SMC | 635-635 | 1 | 2 | 600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 46 | 282 | 25560 (07h06m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
8 | 39 | 8 | 14 | 34 | 7 | 7 | 7 | 34 | 7 | 7 | 7 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
7 | 7 | 7 | 7 | 7 | 7 | 7 | 7 | 6 | 5 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 1 | 1 | 6 | 6 | 6 | 2 | 2 | 2 |
Field13 | Field14 | Field15 | Field16 |
2 | 2 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
19:20 | observing | clear |
20:00 | observing | clear |
21:00 | observing | clear |
22:00 | observing | clear |
23:00 | observing | clear |
00:00 | observing | clear |
01:00 | observing | clear |
02:00 | observing | clear |
03:00 | observing | clear |
04:00 | observing | clear |
05:00 | observing | clear |
06:00 | observing | clear |
06:00 | observing | clear |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day